### v1

# Optical-field-induced dips and splits in nonlinear spectra of selective reflection from high-density atomic vapor

Abstract
We discuss nonlinear spectra of selective reflection from high-density rubidium atomic vapor, where the self-broadening of the resonant transition $5S_{1/2}-5P_{3/2}$ dominates over the Doppler width. In the experiments, the hole-burning technique with probe and pump lasers is used. The reflection of weak probe beam is investigated at four atomic densities in the range $(1.2\text{--}3.6)\cdot 10^{17}$ cm-3 and various pump beam intensities. To enhance the spectral resolution the frequency derivative $\text{d}R/\text{d}\nu$ of the reflection coefficient $R$ recorded and analyzed. The grow of the atomic number density changes the character of self-broadening from inhomogeneous to homogeneous. At the maximal density, the strong pump field split observed spectra on two homogeneously broadened symmetrical resonances. The appearance of the optical field induced resonances can be explained in the frame of ”dressed atomic states” approach. In the range of the lower densities the spectral profiles are in-homogeneously broadened. Spectral profiles of the frequency derivative are separated by the optically saturated dips. The width of such dips is a superposition of the homogeneous part of self-broadening and intensity dependent field broadening. Careful study of the transition from inhomogeneous to homogeneous broadening may be initiate a further development of theory of interatomic interactions in high density atomic gas media.
keywords:
Rubidium atoms, dense atomic gas, self-broadening, Dressed states approach, many-body effects
1 Introduction
In high-density atomic gas, where the dipole self-broadening of resonance transitions is much larger than the Doppler width, the thermal motion of atoms can be neglected Lewis1980 , BoydPRL1991 . The self-broadening $\Gamma_{0}$ in such undisturbed gas of atomic number density $N$ can be evaluated as $\Gamma_{0}=KN$, where $K\sim 10^{-16}$ GHz cm3. The spectral properties of the high-density atomic gases can be recast in the framework of disordered excitons MukamelPRA1994 . Static dipole–dipole interactions between atoms can induce a combination of homogeneous and inhomogeneous broadening of atomic transitions. Possible many-body effects should also be considered. Homogeneous or inhomogeneous character of the line shapes may be tested by applying nonlinear laser spectroscopy methods such as photon echoes or hole burning MukamelJCP1991 . The theoretical publications MukamelPRA1994 , MukamelJCP1991 motivated to conduct an experimental research of the nonlinear selective reflection from the high-density atomic vapor VASPRL1996 . In the experiment with rubidium vapor the probe and pump tunable lasers were used. The probe laser frequency was scanned over the $5S_{1/2}-5P_{3/2}$ transition. The pump laser frequency was fixed in the far wing of the atomic transition, where the absorption length was much longer than the wavelength. The off-resonant optical excitation of rubidium atoms is incoherent due to radiation trapping and interatomic excitation transfer HolsteinPRA1947 , Biberman1947 , VASJQSRT2020 . As a result of a comparison of experimental and calculated reflection spectra a linear relation between self-broadening $\Gamma$ and the ground state population $N_{\text{g}}$ VASPRL1996 is derived $\Gamma=KN_{\text{g}}$.
The observed excitation dependence is explained by static mechanism of the dipole-dipole interactions in the high-density atomic vapor MukamelPRA1994 . Dipole-dipole interactions appeared only between the ground and excited atoms. Subsequent measurements VASPRA1999 , VASJPhysB2009 confirmed the derived linear relation for self-broadening in incoherently excited potassium and rubidium atomic vapors.
The optical resonance saturation of high-density rubidium vapors is investigated in selective reflection experiments using one VASJRLR2021 , BobrovJETPLett2021 , VASJQSRT2022 and two tunable lasers VASNSRJQSRT2024 . Intense laser emission induces a reduction in the width and magnitude of the selective reflection spectral profiles. Field-broadening effects are also discussed.
Interesting nonlinear phenomena may be observed in coherently excited resonant gas, when the Rabi oscillations dominate over relaxation processes Demtroderbook2008 , Boydbook2007 . Observation of the coherent Autler–Townes effect is reported in the seminal publication AutlerTownes1955 . The Autler–Towns components were observed in selective reflection from the high-density potassium vapor VASATPRA2008 . The probe laser frequency was scanned over the $4S_{1/2}-4P_{1/2}$ transition. The pump laser frequency was fixed at the $4S_{1/2}-4P_{3/2}$ transition. The appearance of the Autler–Townes doublet is interpreted within the framework of ”dressed atomic states” approach CohenTannoudjiATbook1996 , CohenTannoudjibook1992 . The measured spectral splitting of the ground state $4S_{1/2}$ is defined by the Rabi frequency.
It has been shown theoretically KaiserPRA2019 that the ”dressed dense atomic gases” may be used to explore the many-body effects. The main goal of our current research is to identify experimental conditions for the formation of the ”dressed states” in the high-density atomic vapors. The optical preparation of new atomic levels can be described by a simple semiclassical model Demtroderbook2008 . In this framework, the optical coherent excitation of an ensemble of two-level atoms with an unperturbed transition frequency $\omega_{ab}$, a transition dipole moment $D_{ab}$ and a homogeneous width $\gamma$ (FWHM) is considered. The excitation rate of atoms, when the frequency of the applied optical field with amplitude $E_{0}$ is tuned at the frequency $\omega_{ab}$, is defined by the ratio of the optically saturated width $\gamma_{\text{sat}}/2$ and the resonance Rabi frequency $\Omega_{\text{R}}$ Demtroderbook2008 :
$\Omega_{\text{R}}=\frac{D_{ab}E_{0}}{\hbar}.$ (1)
When $\Omega_{\text{R}}$ exceeds $\gamma_{\text{sat}}/2$, the Rabi oscillations lead to formation of sidebands at frequencies $\omega_{ab}\pm\Omega_{\text{R}}$. In the absorption spectrum, the symmetrical resonances with a saturated width of $\gamma_{\text{sat}}$ will appear at these frequencies. The frequency interval between the absorption resonances $\Delta\omega_{\text{abs}}$ equals to $2\Omega_{\text{R}}$. According the ”dressed atomic states” approach the strong optical field induces new energy levels by dynamic Stark shifts CohenTannoudjibook1992 , Boydbook2007 .
Figure 1: Optical layout of the setup. Two orthogonally polarized laser beams from pump and probe lasers are combined by polarizing beam splitter (PBS) and focused by lens (L) at internal surface of the cell window. The reflected probe beam is directed by a mirror (M) to a photodiode (PD) through Glan–Thompson polarizer (GT). The frequency of the pump laser is measured by a wavemeter (WM, High-Finesse/Angstrom WS-U). The frequency scan and single-mode regime of the probe laser are monitored by observing saturated absorption in a vapor cell and resonances of reference Fabry–Perot cavity (FPI). The setup also includes the following optical elements: BS–beam splitter, BD–beam dump, QWP–quarter-wave plate, ND–neutral-density filter.
Analysis of selective-reflection spectra is more complicated than analysis of absorption spectra. The spectral profile of unsaturated selective reflection can be roughly approximated by a dispersive curve with a width of $\Gamma_{0}$ BoydPRL1991 , VASPRA1999 . The spectral resolution in this case is limited by slowly decaying wings: $R\sim(\omega-\omega_{ab})$. The derivative $\text{d}R/\text{d}\omega$ gives a bell-shaped spectral profile (absorption-like resonance) with rapidly decreasing wings $R\sim(\omega-\omega_{ab})^{-2}$. Using this technique, the spectral resolution in the selective reflection experiments can be noticeably improved VASATPRA2008 . The resulting spectral curves can be analyzed by using the description of the theoretical absorption spectra Demtroderbook2008 .
In the presented study the derivative of the selective reflection coefficient $\text{d}R/\text{d}\nu$ (frequency $\nu=\omega/2\pi$) are recorded and analyzed at several atomic number densities and varying saturating laser beam intensities.
Figure 2: Spectral dependence of normalized reflection coefficient $(R-R_{0})/R0$ for the probe beam, measured at four number densities: (a) $N_{1}=1.2\times 10^{17}$ cm-3 ($\Gamma_{0}/2\pi=13.2$ GHz), (b) $N_{2}=1.7\times 10^{17}$ cm-3 ($\Gamma_{0}/2\pi=18.7$ GHz), (c) $N_{3}=2.5\times 10^{17}$ cm-3 ($\Gamma_{0}/2\pi=27.5$ GHz), and (d) $N_{4}=3.6\times 10^{17}$ cm-3 ($\Gamma_{0}/2\pi=39.6$ GHz).
Figure 3: Frequency derivatives of the reflection coefficient $\text{d}R/\text{d}\nu$ at four different number densities $N_{1}$, $N_{2}$, $N_{3}$, $N_{4}$ and three selected values of the pump beam intensity $I_{\text{pump}}=0$ kW/cm2 (a), $I_{\text{pump}}=4.8$ kW/cm2 (b), $I_{\text{pump}}=8.8$ kW/cm2 (c). The bottom curves show saturated-absorption spectra of the probe beam for the reference rubidium cell. The frequency scale is the same as in Fig. 2.
2 Experimental setup
Nonlinear selective reflection from the window–vapor interface in the high-temperature vapor cell with natural abundance of 85Rb and 87Rb isotopes was studied at the transition $5S_{1/2}-5P_{3/2}$ (780 nm). The transition consists hyperfine (hfs) components of $^{85}Rb$ and 87Rb isotopes. A brief description of the used vapor cell is given in VASselffoc2019 . The experiments are conducted at four atomic number densities $N_{i}$ in the range $(1.2\text{--}3.6)\cdot 10^{17}$ cm-3 which defined by the temperatures of the cell from $367$ to $427^{\circ}$C. Self-broadening $\Gamma_{0}=KN$ of the working transition is estimated by using the factor $K/2\pi=(1.1\pm 0.17)\cdot 10^{-16}$ GHz cm3 from AdamsJPhysB2011 .
The setup is practically the same as it used in VASNSRJQSRT2024 . Spectroscopic measurements are performed using two tunable diode lasers: a low-power probe laser and a high-power pump laser. The probe lasers frequency is scanned over all hfs-components of $5S_{1/2}-5P_{3/2}$ transition. To calibrate the probe lasers frequency $\nu_{\text{pr}}$, we record the saturated absorption spectrum of the reference rubidium vapor cell with natural abundance of 85Rb and 87Rb isotopes at a room temperature. In our research the hfs-component $5S_{1/2}(F=3)-5P_{3/2}(F^{\prime}=4)$ in 85Rb atoms is chosen as the frequency reference $\nu_{0}$. The pump lasers frequency $\nu_{\text{pump}}$ is fixed at the reference $\nu_{0}$. Schematic of our setup is presented in Fig. 1.
The pump and probe optical beams with orthogonal linear polarization are combined on polarizing beam splitter (PBS) and then focused by lens L onto the inner surface of the cells input window. Application of the pump and probe beams with orthogonal polarizations allows to eliminate possible four-wave mixing at the window/rubidium vapor interface VASFWMPRA1997 , CASFWMQuEl2023 . The pump beam intensity can be varied from zero to 8.8 kW cm-2. The probe beam intensity $I+\text{pr}$ is kept below 4 W cm-2 to avoid a distortion of the recorded curves which could be induced by the probe beam. The probe beam is incident at the 64 mrad. The angle of the incidence for the pump beam is kept at 4 mrad. Probe beam reflected from the interface widow/vapor is directed to a PD photodetector. Scattered radiation from the saturating beam is additionally suppressed by a GT polarizer. The reflection signal from the PD photodetector is recorded using a digital oscilloscope and a computer. The measured selective reflection coefficient $R$ is normalized by the non-resonant reflection coefficient $R_{0}\approx 8.5$% from the window/vacuum interface as $\delta R=(R-R_{0})/R_{0}$. To ensure the higher spectral resolution, we record the derivative of the selective reflection coefficient $\text{d}R/\text{d}\nu$ by computer processing of the experimental data.
3 Experimental results and discussions
The spectral dependencies of the normalized reflection coefficient $\delta R$, recorded at different atomic number densities and pump intensities, are shown in Fig. 2. The frequency scale is expressed as detuning $\Delta_{\text{pr}}/2\pi=(\nu_{\text{pr}}-\nu_{0})$ of the probe laser from the hfs component $5S_{1/2}(F=3)-5P_{3/2}(F^{\prime}=4)$. The pump laser frequency $\nu_{\text{pump}}$ was fixed at the reference frequency $\nu_{0}$. The resonance optical saturation modifies the spectral profiles of $\delta R$, including splitting of the spectra. The width of the unsaturated transition $5S_{1/2}-5P_{3/2}$, which is a combination of self-broadening and hfs-splitting, can be estimated as the frequency interval between minimum and maximum of the $\delta R$ profiles in absence of the pump beam BoydPRL1991 , VASPRA1999 . More reliable results can be reached by measuring the frequency intervals $\Delta\nu_{\text{SR}}$ between ”zeros” of the $\text{d}R/\text{d}\nu$ function VASPRA1999 . The recorded derivatives of the unsaturated reflection coefficient $\text{d}R/\text{d}\nu$ are presented in Fig. 3(a).
According to MukamelPRA1994 in the high-density atomic gas the dipole self-broadening can be a combination of homogeneous and inhomogeneous widths. The hole burning technique can help to estimate contributions of the homogeneous and inhomogeneous parts to the linewidth. In the Figs. 3(b) and 3(c) the nonlinear spectra of frequency derivative $\text{d}R/\text{d}\nu$ are presented. The dips at the pump laser frequency separate the spectra of $\text{d}R/\text{d}\nu$ on two resonances. At densities $N_{1}$, $N_{2}$ and $N_{3}$ the resonances are asymmetrical. The spectral difference between resonances can be attributed to the influence of the hfs-components. Such properties of the spectra indicate that the resonances at densities $N_{1}$, $N_{2}$ and $N_{3}$ are inhomogeneously broadened. Dips are results of the optical saturation of the resonance transitions Demtroderbook2008 . The width of each dip is a superposition of the homogenous part of self-broadening and field broadening Demtroderbook2008 .
Curves in Figs. 3(b), 3(c), which recorded at the maximal density $N_{4}$, demonstrate splitting of spectra of $\text{d}R/\text{d}\nu$ on two symmetrical resonances. It means that the symmetrical resonances are homogeneously broadened. The spectral properties of the symmetrical resonances can be explained by Rabi flopping oscillations in the simple semiclassical model Demtroderbook2008 or by dynamic Stark shifts in the ”dressed states” description CohenTannoudjibook1992 , Boydbook2007 . In both representations the frequency interval between symmetrical resonances is defined by the Rabi frequency. We shall note that a definition of the resonance Rabi frequency $\Omega_{R}$ in Demtroderbook2008 and a definition of the Rabi frequency $\Omega$ in CohenTannoudjibook1992 , Boydbook2007 differ by factor two, as result, $\Omega=2\Omega_{R}$. For the working transition $5S_{1/2}-5P_{3/2}$ in rubidium atoms, the frequency interval $\Delta\nu=2\Omega_{R}$ can be expressed as a function of the pump intensity $I_{\text{pump}}$ VASJRLR2021 :
$\Delta\nu~\text{[GHz]}=2.53\cdot(I_{\text{pump}}\text{[kW/cm}^{-2}\text{]})^{1/2}.$ (2)
The measurements of the splitting interval $\Delta\nu_{\text{SP}}$ between the symmetrical resonances confirm our suggestion, that new levels can be explained by ”atomic dressed states” description, are showed up in our experiment. The experimental data and fitting lines are presented in Fig. 4.
Figure 4: The values of measured splitting $\Delta\nu_{\text{SP}}$ are shown as data points. The lines are results of fit by a linear function $\Delta\nu_{\text{SP}}=a\Delta\nu+b$, where aa–slope of the fitting line, bb–intercept of fitting line with the vertical axis. The inset shows the slope dependence on the atomic number density NN.
In the Table 1 the fitted parameters $a$ and $b$ are presented for the different atomic densities $N_{i}$.
Table 1: Fitted parameters for different atomic densities.
$N_{i}$, $10^{17}$ cm-3 Slope $a$ Intercept $b$, GHz
3.6 $0.93\pm 0.16$ $0.20\pm 0.95$
2.5 $0.59\pm 0.08$ $1.73\pm 0.48$
1.7 $0.42\pm 0.03$ $2.74\pm 0.20$
1.2 $\pm 0.02 3.01±0.13$$\pm 0.1$
The slopes of fitted lines vs density are presented in the inset in Fig. 4. The theoretical slope, obtained by the expression (2), is equal to ”unity”. The difference between the ”unity” and the slope of the fitted line for the maximal density $N_{4}$ is near 7%. This small difference supports our interpretation of the experimental results by ”dressed atomic states” description CohenTannoudjibook1992 , Boydbook2007 . In Table 1 the small values of the slope a at the densities $N_{1}$ and $N_{2}$ reveal the inhomogeneous character of the line-shapes. The frequency intervals $\Delta\nu_{\text{SP}}$ in Fig. 4 are measured between maximums of spectral dependences of derivative $\text{d}R/\text{d}\nu$, separated by the optically saturated dips. The corresponding intercepts $b$ at $N_{1}$ and $N_{2}$ in Table 1 can be associated with homogeneous part of the widths. The intermediate values of $a$ and $b$, measured at the density $N_{3}$, may be explained by a transition from the inhomogeneous self-broadening to the homogeneous self-broadening in this density region. The transition may be a reason for a possible reduction of the contribution of the hfs-components to spectral shapes of the symmetrical resonances in Fig. 3. May be there is an analogy between the reduction of hfs-components in rubidium resonance transition due to the density variation (Fig. 3) and the density dependence of Zeeman splitting of D-lines in potassium vapor VASZeemanPRA1997 . Probably these effects may be associated with many-body interactions in high-density atomic gases.
4 Conclusions
This paper presents the results of an experimental study of nonlinear selective reflection from high-density rubidium atomic vapor, where self-broadening of the resonance transition dominates over Doppler width. In the experiment the pump-probe spectroscopic technique with two lasers are used. The reflection spectra of weak probe beam are recorded at several atomic densities in the range $(1.2\text{--}3.6)\cdot 10^{17}$ cm-3 and various pump beam intensities. With growing density, the broadening character was changed from inhomogeneous to homogeneous. The strong pump field splits the spectra on two resonances. At maximal density, these homogeneously broadened symmetrical resonances are results of the Rabi oscillations Demtroderbook2008 . The appearance of the symmetrical resonances can be explained more strictly in the frame of ”dressed atomic states” approach CohenTannoudjibook1992 , Boydbook2007 . At the maximal density the experimental conditions are appropriated for research of the dressed atomic states. Now we can say that the main goal of our research is reached. At lower densities the spectral profiles are inhomogeneously broadened. Spectral profiles of the frequency derivative are split by the optically saturated dips. At low densities the broadening of spectral profiles is inhomogeneous. Spectral profiles of the frequency derivative are split by the optically saturated dips. The width of such dips is a combination of the homogeneous part of self-broadening and intensity dependent field broadening. Additional studies of the transition from inhomogeneous to homogeneous broadening may initiate a further development of a theory of interatomic interactions in a high-density atomic gas.
In the nanocells SarkisyanPRL2012 , SarkisyanPRL2018 the dressed atomic states can be investigated in absorption spectra of the dense atomic gases. The unique nonlinear properties of ultrathin layers of ”dressed dense atomic gases” may be useful for the development of new quantum devices. The ”dressed states” spectroscopy may be applied for a study of many-body interactions in ultracold nonideal plasmas with the resonance transitions in the visible range ZelenerJETPL2021 , ZelenerPRL2024 .
CRediT authorship contribution statement
Vladimir Sautenkov: Conceptualization, Methodology, Validation, Writing – original draft, Funding acquisition. Sergey Saakyan: Investigation, Software, Validation, Formal analysis, Writing – review & editing. Andrei Bobrov: Formal analysis, Methodology, Writing – review & editing. Boris B. Zelener: Supervision, Resources, Writing – review & editing, Funding acquisition.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
Data will be made available upon request.
Acknowledgements
The research has been supported by the Ministry of Science and Higher Education of the Russian Federation (State Assignment No. 075-00269-25-00).